Find the slope of the line.
3x - 5y = 5

Answer by checkley71(8403)

3X-5Y=5 OR -5Y=-3X+5 OR 5Y=3X-5 OR Y=3/5X-1 THEN THE SLOPE IS 3/5
Y=mX+b WHERE THE m VALUE IS THE SLOPE OF THE LINE.